Assume that you are supplied with a translational potentiometer sensor with the following properties;

* 100 V DC Power Supply,

* 1 resistive wire: 40 mm long with 1mm2 cross-section area. ( resistivity=0,00025 Ohm m)

 It is also given to you that;

 -In the initial state (where no force is applied to the system) of the potentiometer, movable arm of the output voltage positioner covers 90% (length) of the resistive wire.

- When 10N of force is applied to the system you observe 6mm displacement (in (-) y direction) of the movable arm of the output voltage positioner

 

Calculate what is the output voltage change caused by the 10N of force? 


1
Expert's answer
2021-11-05T03:39:49-0400

Determine the maximum resistance:


"R=\rho L/A=10\Omega."

Determine the current:


"I=V/R=10\text A."


The voltage across the arm without load:


"V_1=IR_1=(1-90/100)IR=10\text V."

The voltage across the arm with load:


"V_2=IR_2=IR·\frac{0.1·40+6}{40}=25\text{V}."

The output voltage change:


"\Delta V=V_2-V_1=15\text{ V}."